### v1

# On solutions of reduced biquaternion equality constrained least squares problem and their relative forward error bound 22footnotemark: 2

Abstract
This study focuses on addressing the challenge of solving the reduced biquaternion equality constrained least squares (RBLSE) problem. We develop algebraic techniques to derive both complex and real solutions for the RBLSE problem by utilizing the complex and real forms of reduced biquaternion matrices. Additionally, we conduct a perturbation analysis for the RBLSE problem and establish an upper bound for the relative forward error of these solutions. Numerical examples are presented to illustrate the effectiveness of the proposed approaches and to verify the accuracy of the established upper bound for the relative forward errors.
Keywords. Real representation matrix, Reduced biquaternion, Complex representation matrix, Equality constrained least squares problem.
AMS subject classification. 15A60, 15B33, 65F20, 65H10.
1 Introduction
Reduced biquaternions, first introduced by Segre in $1892$, have become fundamental in various domains such as signal processing, image analysis, and neural networks [4, 5, 6, 7]. Unlike standard quaternions, reduced biquaternions, often referred to as commutative quaternions, possess the unique property of commutative multiplication. This property not only simplifies mathematical operations but also significantly boosts computational efficiency, as demonstrated in [9, 10].
The wide-ranging applications of reduced biquaternions have spurred significant research into both the theoretical and numerical aspects of matrices and matrix equations involving commutative quaternions. For instance, algorithms for calculating eigenvalues, eigenvectors, and singular value decompositions of reduced biquaternion matrices are detailed in [10]. Zhang et al. explored the singular value decomposition and generalized inverse of these matrices in [14], and further investigated the diagonalization process in [15], where they established the necessary and sufficient conditions for diagonalization and introduced two numerical methods to facilitate this task. In [3], authors discussed the LU decomposition of reduced biquaternion matrices. In [8], Guo et al. explored eigen-problems of reduced biquaternion matrices, while in [11], Yuan et al. discussed the Hermitian solution of the reduced biquaternion matrix equation $\left(AXB,CXD\right)=\left(E,F\right)$. In [2], Ding et al. explored the solutions to matrix equations $XF-AX=BY$ and $XF-A\widetilde{X}=BY$ over commutative quaternions.
In many practical applications, determining the solution to linear system, typically expressed as $AX\approx B$, is a common challenge. The least squares method is a well-known approach for addressing this problem. Prior research, such as the work in [12], has explored least squares solutions for matrix equations like $AX=B$ and $AXC=B$ over commutative quaternions. The equality constrained least squares problem is a further study of the least squares problem, and in [13], the authors discussed solution techniques for computing the reduced biquaternion solution of the reduced biquaternion equality constrained least squares (RBLSE) problem. In this manuscript, we advance the research by developing methods to obtain both complex and real solutions for the RBLSE problem. Additionally, we establish an upper bound for the relative forward error associated with these solutions, ensuring the accuracy and reliability of our method in addressing RBLSE problems.
The organization of the manuscript is as follows: Section 2 covers the notation and preliminary concepts. Section 3 outlines the method for finding complex solution of the RBLSE problem, while Section 4 focuses on the technique for finding real solutions to the same problem. Lastly, Section 5 presents the numerical validation of the proposed methods.
2 Notation and preliminaries
2.1 Notation
In this paper, we adopt the following notations: $\mathbb{Q}_{\mathbb{R}}^{m\times n}$, ${\mathbb{C}}^{m\times n}$, and ${\mathbb{R}}^{m\times n}$ represent the sets of all $m\times n$ reduced biquaternion, complex, and real matrices. The notations $\left\lVert\cdot\right\rVert_{2}$ and $\left\lVert\cdot\right\rVert_{F}$ refer to the $2$-norm and Frobenius norm, respectively. For matrices $A\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$ and $B\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $\left[A,B\right]$ represents their concatenation, i.e., $\begin{bmatrix}A&B\end{bmatrix}\in\mathbb{Q}_{\mathbb{R}}^{m\times(n+d)}$. For a matrix $A\in{\mathbb{C}}^{m\times n}$, the symbols $A^{T}$, $A^{H}$, and $A^{\dagger}$ correspond to the transpose, conjugate transpose, and Moore-Penrose inverse of $A$, respectively. Additionally, $A^{-1}$ indicates the inverse of $A$ if it exists.
The Matlab functions $randn(m,n)$ and $rand(m,n)$ produce $m\times n$ matrices filled with normally distributed and uniformly distributed random values, respectively. The abbreviations RBLSE (reduced biquaternion equality constrained least squares) and LSE (equality constrained least squares) will be used throughout this paper.
2.2 Preliminaries
A reduced biquaternion is defined as $\zeta=p_{0}+p_{1}\textit{{i}}+p_{2}\textit{{j}}+p_{3}\textit{{k}}$, which can also be expressed as $\zeta=\left(p_{0}+p_{1}\textit{{i}}\right)+\left(p_{2}+p_{3}\textit{{i}}\right)\textit{{j}}=r_{1}+r_{2}\textit{{j}}$. Here, $p_{i}\in{\mathbb{R}}$ for $i=0,1,2,3$ and $r_{i}\in{\mathbb{C}}$ for $i=1,2$. The defining properties include $\textit{{i}}^{2}=\textit{{k}}^{2}=-1$ and $\textit{{j}}^{2}=1$, with multiplication rules $\textit{{j}}=\textit{{j}}\textit{{i}}=\textit{{k}},\;\textit{{j}}\textit{{k}}=\textit{{k}}\textit{{j}}=\textit{{i}},$ i , and $\textit{{k}}\textit{{i}}=\textit{{i}}\textit{{k}}=-\textit{{j}}$. The norm of $\zeta$ is computed as $\left\lVert\zeta\right\rVert=\sqrt{p_{0}^{2}+p_{1}^{2}+p_{2}^{2}+p_{3}^{2}}=\sqrt{|r_{1}|^{2}+|r_{2}|^{2}}$. For a matrix $M=\left(m_{ij}\right)\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, the Frobenius norm [11] is given by:
$\left\lVert M\right\rVert_{F}=\sqrt{\sum_{i=1}^{m}\sum_{j=1}^{n}\left\lVert m_{ij}\right\rVert^{2}}.$ (1)
Suppose $M=M_{0}+M_{1}\textit{{i}}+M_{2}\textit{{j}}+M_{3}\textit{{k}}=N_{1}+N_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, where $M_{i}\in{\mathbb{R}}^{m\times n}$ for $i=0,1,2,3$ and $N_{i}\in{\mathbb{C}}^{m\times n}$ for $i=1,2$. The real and complex representations of matrix $M$, denoted as $M^{R}$ and $M^{C}$, are given by:
$M^{R}=\begin{bmatrix}M_{0}&-M_{1}&M_{2}&-M_{3}\\ M_{1}&M_{0}&M_{3}&M_{2}\\ M_{2}&-M_{3}&M_{0}&-M_{1}\\ M_{3}&M_{2}&M_{1}&M_{0}\end{bmatrix},\;\;M^{C}=\begin{bmatrix}N_{1}&N_{2}\\ N_{2}&N_{1}\end{bmatrix}.$ (2)
Let $M_{c}^{R}$ represent the first block column of the matrix $M^{R}$. More precisely, $M_{c}^{R}$ is defined as: $M^{R}_{c}=\left[M_{0}^{T},M_{1}^{T},M_{2}^{T},M_{3}^{T}\right]^{T}.$ The matrix $M^{R}$ can be represented using $M^{R}_{c}$ as:
$M^{R}=\left[M_{c}^{R},Q_{m}M_{c}^{R},R_{m}M_{c}^{R},S_{m}M_{c}^{R}\right],$ (3)
where
$Q_{m}=\begin{bmatrix}0&-I_{m}&0&0\\ I_{m}&0&0&0\\ 0&0&0&-I_{m}\\ 0&0&I_{m}&0\end{bmatrix},\;R_{m}=\begin{bmatrix}0&0&I_{m}&0\\ 0&0&0&I_{m}\\ I_{m}&0&0&0\\ 0&I_{m}&0&0\end{bmatrix},\mbox{and}\;S_{m}=\begin{bmatrix}0&0&0&-I_{m}\\ 0&0&I_{m}&0\\ 0&-I_{m}&0&0\\ I_{m}&0&0&0\end{bmatrix}.$ (4)
We have
$\left\lVert M\right\rVert_{F}=\frac{1}{2}\left\lVert M^{R}\right\rVert_{F}=\left\lVert M^{R}_{c}\right\rVert_{F}.$ (5)
Lemma 2.1.
Let $P\in{\mathbb{R}}^{m\times n}$, $Q\in{\mathbb{R}}^{m\times d}$, $R\in{\mathbb{R}}^{m\times p}$, and $S\in{\mathbb{R}}^{m\times q}$. If $\|P\|_{F}=\|Q\|_{F}=\|R\|_{F}=\|S\|_{F}$, then we have
$\left\lVert P\right\rVert_{F}=\frac{1}{2}\left\lVert\left[P,Q,R,S\right]\right\rVert_{F}.$
Let $M_{c}^{C}$ represent the first block column of the matrix $M^{C}$. More precisely, $M_{c}^{C}$ is defined as: $M^{C}_{c}=\left[M_{1}^{T},M_{2}^{T}\right]^{T}.$ The matrix $M^{C}$ can be represented using $M^{C}_{c}$ as:
$M^{C}=\left[M_{c}^{C},P_{m}M_{c}^{C}\right],$ (6)
where
$P_{m}=\begin{bmatrix}0&I_{m}\\ I_{m}&0\end{bmatrix}.$ (7)
We have
$\left\lVert M\right\rVert_{F}=\frac{1}{\sqrt{2}}\left\lVert M^{C}\right\rVert_{F}=\left\lVert M^{C}_{c}\right\rVert_{F}.$ (8)
Lemma 2.2.
Let $P\in{\mathbb{C}}^{m\times n}$ and $Q\in{\mathbb{C}}^{m\times d}$. If $\|P\|_{F}=\|Q\|_{F}$, then we have
$\left\lVert P\right\rVert_{F}=\frac{1}{\sqrt{2}}\left\lVert\left[P,Q\right]\right\rVert_{F}.$
Lemma 2.3.
For $\alpha\in{\mathbb{R}}$, $\beta\in{\mathbb{C}}$, $P,Q\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, and $R\in\mathbb{Q}_{\mathbb{R}}^{n\times t}$, the following relationships hold:
  1. $(a)$
    $P=Q\iff P^{C}=Q^{C}\iff P^{R}=Q^{R}$.
  2. $(b)$
    $\left(P+Q\right)^{R}=P^{R}+Q^{R}$, $\left(P+Q\right)^{C}=P^{C}+Q^{C}$.
  3. $(c)$
    $\left(\alpha P\right)^{R}=\alpha P^{R}$, $\left(\beta P\right)^{C}=\beta P^{C}$.
  4. $(d)$
    $\left(PR\right)^{R}=P^{R}R^{R}$, $\left(PR\right)^{C}=P^{C}R^{C}$.
3 An algebraic technique for complex solution of RBLSE problem
This section focuses on an algebraic approach to derive the complex solution for the RBLSE problem. The method is based on analyzing the solution of the associated complex LSE problem. Suppose
$A$ $=$ $M_{1}+M_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n},\;\;\;\;B=N_{1}+N_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d},$ (9)
$C$ $=$ $R_{1}+R_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n},\;\;\;\;D=S_{1}+S_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ (10)
We will limit our discussion to the scenario where $m\geq n+d$, and both matrices $C$ and $C^{C}_{c}$ have full row rank. The RBLSE problem can then be stated as follows:
$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert AX-B\right\rVert_{F}\;\;\;\textrm{subject to}\;\;\;CX=D.$ (11)
Consider a complex LSE problem
$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}\;\;\;\textrm{subject to}\;\;\;C^{C}_{c}X=D^{C}_{c}.$ (12)
Let the QR factorization of $\left(C^{C}_{c}\right)^{H}$ be as follows:
$\left(C^{C}_{c}\right)^{H}=\widetilde{Q}\begin{bmatrix}\widetilde{R}\\ 0\end{bmatrix},$ (13)
where $\widetilde{Q}\in{\mathbb{C}}^{n\times n}$ is a unitary matrix and $\widetilde{R}\in{\mathbb{C}}^{2p\times 2p}$ is a nonsingular upper triangular matrix. Now, partition $A^{C}_{c}\widetilde{Q}$ as:
$A^{C}_{c}\widetilde{Q}=\left[\widetilde{P}_{1},\widetilde{P}_{2}\right],$ (14)
where $\widetilde{P}_{1}\in{\mathbb{C}}^{2m\times 2p}$ and $\widetilde{P}_{2}\in{\mathbb{C}}^{2m\times(n-2p)}$.
Theorem 3.1.
Consider the RBLSE problem outlined in (11) and the complex LSE problem in (12), with notations specified in (13) and (14). For a matrix $X\in{\mathbb{C}}^{n\times d}$, $X$ is a complex solution of the RBLSE problem (11) if and only if $X$ solves the complex LSE problem (12). In this scenario, the unique solution $X$ with minimum norm can be expressed as:
$X=\widetilde{Q}\begin{bmatrix}\left(\widetilde{R}^{H}\right)^{-1}D^{C}_{c}\\ \widetilde{P}_{2}^{\dagger}\left(B^{C}_{c}-\widetilde{P}_{1}\left(\widetilde{R}^{H}\right)^{-1}D^{C}_{c}\right)\end{bmatrix}.$ (15)
Proof.
If $X\in{\mathbb{C}}^{n\times d}$ is a solution of the complex LSE problem (12), then
$\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}=\textrm{min},\;\;C^{C}_{c}X=D^{C}_{c}.$ (16)
The Frobenius norm of a complex matrix remains invariant under unitary transformations. Since the matrix $P_{m}$ in (7) is unitary, it follows that:
$\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}=\left\lVert P_{m}\left(A^{C}_{c}X-B^{C}_{c}\right)\right\rVert_{F}.$
Using equations (2), (6), (8), along with Lemmas 2.2 and 2.3, we obtain
$\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}$ $=$ $\frac{1}{\sqrt{2}}\left\lVert\left[\left(A^{C}_{c}X-B^{C}_{c}\right),P_{m}\left(A^{C}_{c}X-B^{C}_{c}\right)\right]\right\rVert_{F}$
$=$ $\frac{1}{\sqrt{2}}\left\lVert\left[A^{C}_{c}X,P_{m}A^{C}_{c}X\right]-\left[B^{C}_{c},P_{m}B^{C}_{c}\right]\right\rVert_{F}$
$=$ $\frac{1}{\sqrt{2}}\left\lVert\left[A^{C}_{c},P_{m}A^{C}_{c}\right]\begin{bmatrix}X&0\\ 0&X\end{bmatrix}-\left[B^{C}_{c},P_{m}B^{C}_{c}\right]\right\rVert_{F}$
$=$ $\frac{1}{\sqrt{2}}\left\lVert A^{C}X^{C}-B^{C}\right\rVert_{F}$
$=$ $\frac{1}{\sqrt{2}}\left\lVert\left(AX-B\right)^{C}\right\rVert_{F}$
$=$ $\left\lVert AX-B\right\rVert_{F}.$
From (16), we obtain
$\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}=\left\lVert AX-B\right\rVert_{F}=\textrm{min},$ (17)
and
$\left[C^{C}_{c},P_{p}C^{C}_{c}\right]\begin{bmatrix}X&0\\ 0&X\end{bmatrix}=\left[D^{C}_{c},P_{p}D^{C}_{c}\right].$
Using (6), we know that $C^{C}=\left[C^{C}_{c},P_{p}C^{C}_{c}\right]$ and $D^{C}=\left[D^{C}_{c},P_{p}D^{C}_{c}\right]$. Applying this, we get
$C^{C}X^{C}$ $=$ $D^{C},$
$\left(CX\right)^{C}$ $=$ $D^{C},$
$CX$ $=$ $D.$ (18)
By combining (17) and (18), we conclude that $X\in{\mathbb{C}}^{n\times d}$ is a complex solution to the RBLSE problem (11), and vice versa.
To find the expression for $X$, we solve the complex LSE problem (12). Set
$\widetilde{Q}^{H}X=\begin{bmatrix}Y\\ Z\end{bmatrix},$
where $Y\in{\mathbb{C}}^{2p\times d}$ and $Z\in{\mathbb{C}}^{(n-2p)\times d}$. Equation 12 can be rewritten as
$\min_{X}\left\lVert A^{C}_{c}\widetilde{Q}\widetilde{Q}^{H}X-B^{C}_{c}\right\rVert_{F}\;\;\;\textrm{subject to}\;\;\;C^{C}_{c}\widetilde{Q}\widetilde{Q}^{H}X=D^{C}_{c}.$ (19)
Utilizing (13), we have
$C^{C}_{c}\widetilde{Q}\widetilde{Q}^{H}X=D^{C}_{c}\implies\begin{bmatrix}\widetilde{R}^{H}&0\end{bmatrix}\begin{bmatrix}Y\\ Z\end{bmatrix}=D^{C}_{c}.$
Since $\widetilde{R}^{H}$ is a nonsingular matrix, we get $Y=\left(\widetilde{R}^{H}\right)^{-1}D^{C}_{c}$. Using (14), Equation 19 takes the form
$\min_{Z}\left\lVert\widetilde{P}_{2}Z-\left(B^{C}_{c}-\widetilde{P}_{1}Y\right)\right\rVert_{F}.$
The minimum norm solution of the above least squares problem is $Z=\widetilde{P}_{2}^{\dagger}\left(B^{C}_{c}-\widetilde{P}_{1}Y\right)$. Thus, we can derive the desired expression for $X$. ∎
Next, we aim to examine how perturbations in $A$, $B$, $C$, and $D$ affect the complex solution $X_{CL}$ of the RBLSE problem (11). Let
$\widehat{A}=A+\Delta A,\;\widehat{B}=B+\Delta B,\;\widehat{C}=C+\Delta C,\;\textrm{and}\;\widehat{D}=D+\Delta D,$ (20)
where $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ represent the perturbations of the input data $A$, $B$, $C$, and $D$, respectively. We assume that the perturbation $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ are small enough to guarantee that the perturbed matrices $\widehat{C}$ and $\widehat{C}^{C}_{c}$ retain full row rank. These perturbations are measured normwise by the smallest $\epsilon$ for which
$\left\lVert\Delta A\right\rVert_{F}\leq\epsilon\left\lVert A\right\rVert_{F},\;\left\lVert\Delta B\right\rVert_{F}\leq\epsilon\left\lVert B\right\rVert_{F},\;\left\lVert\Delta C\right\rVert_{F}\leq\epsilon\left\lVert C\right\rVert_{F},\;\textrm{and}\;\left\lVert\Delta D\right\rVert_{F}\leq\epsilon\left\lVert D\right\rVert_{F}.$ (21)
Let $\widehat{X}_{CL}$ be the complex solution to the perturbed RBLSE problem
$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert\widehat{A}X-\widehat{B}\right\rVert_{F}\;\;\textrm{subject to}\;\;\widehat{C}X=\widehat{D},$ (22)
and let $\Delta X_{CL}=\widehat{X}_{CL}-X_{CL}$.
Theorem 3.2.
Consider the RBLSE problem outlined in (11) and the perturbed RBLSE problem described in (22). If the perturbations $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ are sufficiently small, as described in (21), then we have
$\begin{split}\frac{\left\lVert\Delta X_{CL}\right\rVert_{F}}{\left\lVert X_{CL}\right\rVert_{F}}\leq\epsilon&\left(\mathcal{K}_{A}^{C}\left(\frac{\left\lVert D^{C}_{c}\right\rVert_{F}}{\left\lVert C^{C}_{c}\right\rVert_{F}\left\lVert X_{CL}\right\rVert_{F}}+1\right)+\mathcal{K}_{B}^{C}\left(\frac{\left\lVert B^{C}_{c}\right\rVert_{F}}{\left\lVert A^{C}_{c}\right\rVert_{F}\left\lVert X_{CL}\right\rVert_{F}}+1\right)\right.\\ &\left.+\left(\mathcal{K}_{B}^{C}\right)^{2}\left(\frac{\left\lVert C^{C}_{c}\right\rVert_{F}}{\left\lVert A^{C}_{c}\right\rVert_{F}}\left\lVert A^{C}_{c}\mathcal{L}_{c}\right\rVert_{2}+1\right)\frac{\left\lVert R_{c}\right\rVert_{F}}{\left\lVert A^{C}_{c}\right\rVert_{F}\left\lVert X_{CL}\right\rVert_{F}}\right)+O(\epsilon^{2})\equiv U_{CL},\end{split}$ (23)
where
$\mathcal{K}_{B}^{C}$ $=$ $\left\lVert A^{C}_{c}\right\rVert_{F}\left\lVert\left(A^{C}_{c}P_{c}\right)^{\dagger}\right\rVert_{2},\;\;\;\mathcal{K}_{A}^{C}=\left\lVert C^{C}_{c}\right\rVert_{F}\left\lVert\mathcal{L}_{c}\right\rVert_{2},\;\;\;\mathcal{L}_{c}=\left(I_{n}-\left(A^{C}_{c}P_{c}\right)^{\dagger}A^{C}_{c}\right)\left(C^{C}_{c}\right)^{\dagger},\;\;\;\;$
$P_{c}$ $=$ $I_{n}-\left(C^{C}_{c}\right)^{\dagger}C^{C}_{c},\;\;\;\;\;\;\;\;\;\;\;R_{c}=B^{C}_{c}-A^{C}_{c}X_{CL}.$
Proof.
The perturbed complex LSE problem corresponding to the perturbed RBLSE problem (22) is given by:
$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert\left(\widehat{A}\right)^{C}_{c}X-\left(\widehat{B}\right)^{C}_{c}\right\rVert_{F}\;\;\textrm{subject to}\;\;\left(\widehat{C}\right)^{C}_{c}X=\left(\widehat{D}\right)^{C}_{c}.$ (24)
Using Theorem 3.1, we know that $\widehat{X}_{CL}$ is the solution to the perturbed complex LSE problem (24). From (20) and utilizing Lemma 2.3, we have
$\left(\widehat{A}\right)^{C}_{c}=A^{C}_{c}+\left(\Delta A\right)^{C}_{c},\;\left(\widehat{B}\right)^{C}_{c}=B^{C}_{c}+\left(\Delta B\right)^{C}_{c},\;\left(\widehat{C}\right)^{C}_{c}=C^{C}_{c}+\left(\Delta C\right)^{C}_{c},\;\textrm{and}\;\left(\widehat{D}\right)^{C}_{c}=D^{C}_{c}+\left(\Delta D\right)^{C}_{c}.$
Thus, the perturbed complex LSE problem (24) can be rewritten as:
$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert\left(A^{C}_{c}+\left(\Delta A\right)^{C}_{c}\right)X-\left(B^{C}_{c}+\left(\Delta B\right)^{C}_{c}\right)\right\rVert_{F}\;\;\textrm{subject to}\;\;\left(C^{C}_{c}+\left(\Delta C\right)^{C}_{c}\right)X=\left(D^{C}_{c}+\left(\Delta D\right)^{C}_{c}\right).$ (25)
Using (8) and (21), we can establish the following bounds for the perturbation:
$\left\lVert\left(\Delta A\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert A^{C}_{c}\right\rVert_{F},\;\left\lVert\left(\Delta B\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert B^{C}_{c}\right\rVert_{F},\;\left\lVert\left(\Delta C\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert C^{C}_{c}\right\rVert_{F},\;\textrm{and}\;\left\lVert\left(\Delta D\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert D^{C}_{c}\right\rVert_{F}.$ (26)
With the perturbed problem (25) and the bounds in (26), and using Theorem 3.1, the sensitivity analysis of the complex solution to the RBLSE problem (11) reduces to evaluating the sensitivity of the solution to the complex LSE problem (12). Consequently, the upper bound $U_{CL}$ for the relative forward error of the complex solution to the RBLSE problem can be obtained from [1, Equation $4.11$]. ∎
4 An algebraic technique for real solution of RBLSE problem
This section focuses on an algebraic approach to derive the real solution for the RBLSE problem. The method is based on analyzing the solution of the associated real LSE problem. Suppose
$A$ $=$ $A_{0}+A_{1}\textit{{i}}+A_{2}\textit{{j}}+A_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n},\;\;\;\;B=B_{0}+B_{1}\textit{{i}}+B_{2}\textit{{j}}+B_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d},$ (27)
$C$ $=$ $C_{0}+C_{1}\textit{{i}}+C_{2}\textit{{j}}+C_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n},\;\;\;\;\;D=D_{0}+D_{1}\textit{{i}}+D_{2}\textit{{j}}+D_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ (28)
We will limit our discussion to the scenario where $m\geq n+d$, and both matrices $C$ and $C^{R}_{c}$ have full row rank. The RBLSE problem can then be stated as follows:
$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert AX-B\right\rVert_{F}\;\;\;\textrm{subject to}\;\;\;CX=D.$ (29)
Consider a real LSE problem
$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert A^{R}_{c}X-B^{R}_{c}\right\rVert_{F}\;\;\;\textrm{subject to}\;\;\;C^{R}_{c}X=D^{R}_{c}.$ (30)
Let the QR factorization of $\left(C^{R}_{c}\right)^{T}$ be as follows:
$\left(C^{R}_{c}\right)^{T}=\bar{Q}\begin{bmatrix}\bar{R}\\ 0\end{bmatrix},$ (31)
where $\bar{Q}\in{\mathbb{R}}^{n\times n}$ is an orthonormal matrix and $\bar{R}\in{\mathbb{R}}^{4p\times 4p}$ is a nonsingular upper triangular matrix. Now, partition $A^{R}_{c}\bar{Q}$ as:
$A^{R}_{c}\bar{Q}=\left[\bar{P}_{1},\bar{P}_{2}\right],$ (32)
where $\bar{P}_{1}\in{\mathbb{R}}^{4m\times 4p}$ and $\bar{P}_{2}\in{\mathbb{R}}^{4m\times(n-4p)}$.
Theorem 4.1.
Consider the RBLSE problem outlined in (29) and the real LSE problem in (30), with notations specified in (31) and (32). For a matrix $X\in{\mathbb{R}}^{n\times d}$, $X$ is a real solution of the RBLSE problem (29) if and only if $X$ solves the real LSE problem (30). In this scenario, the unique solution with minimum norm $X$ can be expressed as:
$X=\bar{Q}\begin{bmatrix}\left(\bar{R}^{T}\right)^{-1}D^{R}_{c}\\ \bar{P}_{2}^{\dagger}\left(B^{R}_{c}-\bar{P}_{1}\left(\bar{R}^{T}\right)^{-1}D^{R}_{c}\right)\end{bmatrix}.$ (33)
Proof.
If $X\in{\mathbb{R}}^{n\times d}$ is a solution of the real LSE problem (30), then
$\left\lVert A^{R}_{c}X-B^{R}_{c}\right\rVert_{F}=\textrm{min},\;\;C^{R}_{c}X=D^{R}_{c}.$ (34)
The Frobenius norm of a real matrix remains invariant under orthogonal transformations. Since the matrices $Q_{m}$, $R_{m}$, and $S_{m}$ in (4) are orthogonal, it follows that:
$\left\lVert A^{R}_{c}X-B^{R}_{c}\right\rVert_{F}=\left\lVert Q_{m}\left(A^{R}_{c}X-B^{R}_{c}\right)\right\rVert_{F}=\left\lVert R_{m}\left(A^{R}_{c}X-B^{R}_{c}\right)\right\rVert_{F}=\left\lVert S_{m}\left(A^{R}_{c}X-B^{R}_{c}\right)\right\rVert_{F}.$
Using equations (2), (3), (5), along with Lemmas 2.1 and 2.3, we obtain
$\left\lVert A^{R}_{c}X-B^{R}_{c}\right\rVert_{F}$ $=$ $\frac{1}{2}\left\lVert\left[\left(A^{R}_{c}X-B^{R}_{c}\right),Q_{m}\left(A^{R}_{c}X-B^{R}_{c}\right),R_{m}\left(A^{R}_{c}X-B^{R}_{c}\right),S_{m}\left(A^{R}_{c}X-B^{R}_{c}\right)\right]\right\rVert_{F}$
$=$ $\frac{1}{2}\left\lVert\left[A^{R}_{c}X,Q_{m}A^{R}_{c}X,R_{m}A^{R}_{c}X,S_{m}A^{R}_{c}X\right]-\left[B^{R}_{c},Q_{m}B^{R}_{c},R_{m}B^{R}_{c},S_{m}B^{R}_{c}\right]\right\rVert_{F}$
$=$ $\frac{1}{2}\left\lVert\left[A^{R}_{c},Q_{m}A^{R}_{c},R_{m}A^{R}_{c},S_{m}A^{R}_{c}\right]\begin{bmatrix}X&0&0&0\\ 0&X&0&0\\ 0&0&X&0\\ 0&0&0&X\end{bmatrix}-\left[B^{R}_{c},Q_{m}B^{R}_{c},R_{m}B^{R}_{c},S_{m}B^{R}_{c}\right]\right\rVert_{F}$
$=$ $\frac{1}{2}\left\lVert A^{R}X^{R}-B^{R}\right\rVert_{F}$
$=$ $\frac{1}{2}\left\lVert\left(AX-B\right)^{R}\right\rVert_{F}$
$=$ $\left\lVert AX-B\right\rVert_{F}.$
From (34), we obtain
$\left\lVert A^{R}_{c}X-B^{R}_{c}\right\rVert_{F}=\left\lVert AX-B\right\rVert_{F}=\textrm{min},$ (35)
and
$\left[C^{R}_{c},Q_{p}C^{R}_{c},R_{p}C^{R}_{c},S_{p}C^{R}_{c}\right]\begin{bmatrix}X&0&0&0\\ 0&X&0&0\\ 0&0&X&0\\ 0&0&0&X\end{bmatrix}=\left[D^{R}_{c},Q_{p}D^{R}_{c},R_{p}D^{R}_{c},S_{p}D^{R}_{c}\right].$
Using (3), we know that $C^{R}=\left[C^{R}_{c},Q_{p}C^{R}_{c},R_{p}C^{R}_{c},S_{p}C^{R}_{c}\right]$ and $D^{R}=\left[D^{R}_{c},Q_{p}D^{R}_{c},R_{p}D^{R}_{c},S_{p}D^{R}_{c}\right]$. Applying this, we get
$C^{R}X^{R}$ $=$ $D^{R},$
$(CX)^{R}$ $=$ $D^{R},$
$CX$ $=$ $D.$ (36)
By combining (35) and (36), we conclude that $X\in{\mathbb{R}}^{n\times d}$ is a real solution to the RBLSE problem (29), and vice versa. The expression for $X$ can be obtained by following the proof method of Theorem 3.1. ∎
Next, we aim to examine how perturbations in $A$, $B$, $C$, and $D$ affect the real solution $X_{RL}$ of the RBLSE problem (29). Let
$\widehat{A}=A+\Delta A,\;\widehat{B}=B+\Delta B,\;\widehat{C}=C+\Delta C,\;\textrm{and}\;\widehat{D}=D+\Delta D,$ (37)
where $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ represent the perturbations of the input data $A$, $B$, $C$, and $D$, respectively. We assume that the perturbation $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ are small enough to ensure that the perturbed matrices $\widehat{C}$ and $\widehat{C}^{R}_{c}$ retain full row rank. These perturbations are measured normwise by the smallest $\epsilon$ for which
$\left\lVert\Delta A\right\rVert_{F}\leq\epsilon\left\lVert A\right\rVert_{F},\;\left\lVert\Delta B\right\rVert_{F}\leq\epsilon\left\lVert B\right\rVert_{F},\;\left\lVert\Delta B\right\rVert_{F}\leq\epsilon\left\lVert B\right\rVert_{F},\;\textrm{and}\;\left\lVert\Delta D\right\rVert_{F}\leq\epsilon\left\lVert D\right\rVert_{F}.$ (38)
Let $\widehat{X}_{RL}$ be the real solution to the perturbed RBLSE problem
$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert\widehat{A}X-\widehat{B}\right\rVert_{F}\;\;\textrm{subject to}\;\;\widehat{C}X=\widehat{D},$ (39)
and let $\Delta X_{RL}=\widehat{X}_{RL}-X_{RL}$.
Theorem 4.2.
Consider the RBLSE problem outlined in (29) and the perturbed RBLSE problem described in (39). If the perturbations $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ are sufficiently small, as described in (38), then we have
$\begin{split}\frac{\left\lVert\Delta X_{RL}\right\rVert_{F}}{\left\lVert X_{RL}\right\rVert_{F}}\leq\epsilon&\left(\mathcal{K}_{A}^{R}\left(\frac{\left\lVert D^{R}_{c}\right\rVert_{F}}{\left\lVert C^{R}_{c}\right\rVert_{F}\left\lVert X_{RL}\right\rVert_{F}}+1\right)+\mathcal{K}_{B}^{R}\left(\frac{\left\lVert B^{R}_{c}\right\rVert_{F}}{\left\lVert A^{R}_{c}\right\rVert_{F}\left\lVert X_{RL}\right\rVert_{F}}+1\right)\right.\\ &\left.+\left(\mathcal{K}_{B}^{R}\right)^{2}\left(\frac{\left\lVert C^{R}_{c}\right\rVert_{F}}{\left\lVert A^{R}_{c}\right\rVert_{F}}\left\lVert A^{R}_{c}\mathcal{L}_{r}\right\rVert_{2}+1\right)\frac{\left\lVert R_{r}\right\rVert_{F}}{\left\lVert A^{R}_{c}\right\rVert_{F}\left\lVert X_{RL}\right\rVert_{F}}\right)+O(\epsilon^{2})\equiv U_{RL},\end{split}$ (40)
where
$\mathcal{K}_{B}^{R}$ $=$ $\left\lVert A^{R}_{c}\right\rVert_{F}\left\lVert\left(A^{R}_{c}P_{r}\right)^{\dagger}\right\rVert_{2},\;\;\;\mathcal{K}_{A}^{R}=\left\lVert C^{R}_{c}\right\rVert_{F}\left\lVert\mathcal{L}_{r}\right\rVert_{2},\;\;\;\mathcal{L}_{r}=\left(I_{n}-\left(A^{R}_{c}P_{r}\right)^{\dagger}A^{R}_{c}\right)\left(C^{R}_{c}\right)^{\dagger},\;\;\;\;$
$P_{r}$ $=$ $I_{n}-\left(C^{R}_{c}\right)^{\dagger}C^{R}_{c},\;\;\;\;\;\;\;\;\;\;\;R_{r}=B^{R}_{c}-A^{R}_{c}X_{RL}.$
Proof.
The perturbed real LSE problem corresponding to the perturbed RBLSE problem (39) is given by:
$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert\left(\widehat{A}\right)^{R}_{c}X-\left(\widehat{B}\right)^{R}_{c}\right\rVert_{F}\;\;\textrm{subject to}\;\;\left(\widehat{C}\right)^{R}_{c}X=\left(\widehat{D}\right)^{R}_{c}.$ (41)
Using Theorem 4.1, we know that $\widehat{X}_{RL}$ is the solution to the perturbed real LSE problem (41). From (37) and utilizing Lemma 2.3, we have
$\left(\widehat{A}\right)^{R}_{c}=A^{R}_{c}+\left(\Delta A\right)^{R}_{c},\;\left(\widehat{B}\right)^{R}_{c}=B^{R}_{c}+\left(\Delta B\right)^{R}_{c},\;\left(\widehat{C}\right)^{R}_{c}=C^{R}_{c}+\left(\Delta C\right)^{R}_{c},\;\textrm{and}\;\left(\widehat{D}\right)^{R}_{c}=D^{R}_{c}+\left(\Delta D\right)^{R}_{c}.$
Thus, the perturbed real LSE problem (41) can be rewritten as:
$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert\left(A^{R}_{c}+\left(\Delta A\right)^{R}_{c}\right)X-\left(B^{R}_{c}+\left(\Delta B\right)^{R}_{c}\right)\right\rVert_{F}\;\;\textrm{subject to}\;\;\left(C^{R}_{c}+\left(\Delta C\right)^{R}_{c}\right)X=\left(D^{R}_{c}+\left(\Delta D\right)^{R}_{c}\right).$ (42)
Using (5) and (38), we can establish the following bounds for the perturbation:
$\left\lVert\left(\Delta A\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert A^{R}_{c}\right\rVert_{F},\;\left\lVert\left(\Delta B\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert B^{R}_{c}\right\rVert_{F},\;\left\lVert\left(\Delta C\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert C^{R}_{c}\right\rVert_{F},\;\textrm{and}\;\left\lVert\left(\Delta D\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert D^{R}_{c}\right\rVert_{F}.$ (43)
With the perturbed problem (42) and the bounds in (43), and using Theorem 4.1, the sensitivity analysis of the real solution to the RBLSE problem (29) reduces to evaluating the sensitivity of the solution to the real LSE problem (30). Consequently, the upper bound $U_{RL}$ for the relative forward error of the real solution to the RBLSE problem can be obtained from [1, Equation $4.11$]. ∎
5 Numerical Verification
Building on the previous discussions, this section presents numerical algorithms designed to find special solutions to the RBLSE problem. We also include numerical examples to validate these algorithms. Additionally, we assess the upper bound for the relative forward error of both complex and real solutions to the RBLSE problem.
Algorithm 1 For the complex solution of the RBLSE problem (11)
Input: $A=M_{1}+M_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, $B=N_{1}+N_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $C=R_{1}+R_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n}$, and $D=S_{1}+S_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ We assume $m\geq n+d$ and that both matrices $C$ and $C^{C}_{c}$ have full row rank.
Output: $X$.
  1. 1.
    Find the QR factorization of $\left(C^{C}_{c}\right)^{H}$, as described in (13).
  2. 2.
    Partition the matrix $A^{C}_{c}\widetilde{Q}$ as shown in (14).
  3. 3.
    The complex solution $X$ to the RBLSE problem (11) is given by (15).
Algorithm 2 For the real solution of the RBLSE problem (29)
Input: $A=A_{0}+A_{1}\textit{{i}}+A_{2}\textit{{j}}+A_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, $B=B_{0}+B_{1}\textit{{i}}+B_{2}\textit{{j}}+B_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $C=C_{0}+C_{1}\textit{{i}}+C_{2}\textit{{j}}+C_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n}$, and $D=D_{0}+D_{1}\textit{{i}}+D_{2}\textit{{j}}+D_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ We assume $m\geq n+d$ and that both matrices $C$ and $C^{R}_{c}$ have full row rank.
Output: $X$.
  1. 1.
    Find the QR factorization of $\left(C^{R}_{c}\right)^{T}$, as described in (31).
  2. 2.
    Partition the matrix $A^{R}_{c}\bar{Q}$ as shown in (32).
  3. 3.
    The real solution $X$ to the RBLSE problem (29) is given by (33).
Next, we provide examples to evaluate the effectiveness of the proposed algorithms. All computations were performed using a computer equipped with an Intel Core $i7-9700$ processor at $3.00$ GHz and $16$ GB of RAM, running MATLAB $R2021b$ software.
Example 5.1.
Let $A=M_{1}+M_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, $B=N_{1}+N_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $C=R_{1}+R_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n}$, and $D=S_{1}+S_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}$. Let
$M_{i}$ $=$ $rand(m,n)+rand(m,n)\textit{{i}}\in{\mathbb{C}}^{m\times n},\;\textrm{for}\;i=1,2,$
$N_{i}$ $=$ $rand(m,d)+rand(m,d)\textit{{i}}\in{\mathbb{C}}^{m\times d},\;\textrm{for}\;i=1,2,$
$R_{i}$ $=$ $rand(p,n)+rand(p,n)\textit{{i}}\in{\mathbb{C}}^{p\times n},\;\textrm{for}\;i=1,2,$
$S_{i}$ $=$ $rand(p,d)+rand(p,d)\textit{{i}}\in{\mathbb{C}}^{p\times d},\;\textrm{for}\;i=1,2.$
Take $m=40t$, $n=6t$, $p=2t$, and $d=3$. Here, $t$ is an arbitrary number. We apply Algorithm 1 to determine the complex solution of the RBLSE problem. Let $X_{CL}$ be the complex solution of the RBLSE problem
$\min_{X,R_{c}}\left\lVert R_{c}\right\rVert_{F}\;\;\textrm{subject to}\;\;AX=B+R_{c},\;CX=D.$
Let $\epsilon_{1}=\left\lVert AX_{CL}-(B+R_{c})\right\rVert_{F}$ and $\epsilon_{2}=\left\lVert CX_{CL}-D\right\rVert_{F}$. In Table 1, we compute $\epsilon_{1}$ and $\epsilon_{2}$ for different values of $t$.
$t$ $\epsilon_{1}$ $\epsilon_{2}$
$1$ $8.5131\times 10^{-16}$ $2.0907\times 10^{-15}$
$3$ $7.0956\times 10^{-16}$ $2.5624\times 10^{-15}$
$5$ $1.0499\times 10^{-15}$ $3.7683\times 10^{-15}$
$7$ $1.2804\times 10^{-15}$ $4.5681\times 10^{-15}$
$9$ $1.4232\times 10^{-15}$ $8.1546\times 10^{-15}$
Table 1: Computational accuracy of Algorithm 1 for computing the complex solution of the RBLSE problem (11)
Table 1 shows that the errors $\epsilon_{1}$ and $\epsilon_{2}$ across different values of $t$ are consistently below $10^{-15}$. This indicates that Algorithm 1 is highly effective in determining the complex solution for the RBLSE problem (11).
Example 5.2.
In Example 5.1, we introduce random perturbations $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ to the input matrices $A$, $B$, $C$, and $D$, respectively, to examine how these small perturbations affect the complex solution $X_{CL}$ of the RBLSE problem (11). These perturbations are measured normwise by the smallest $\epsilon$, as in (21). We define the exact relative forward errors for these solutions as $\frac{\|\widehat{X}_{CL}-X_{CL}\|_{F}}{\|X_{CL}\|_{F}}$. Table 2 presents the calculated exact relative forward errors and the corresponding upper bound $U_{CL}$ (calculated using Equation 23) for these solutions across different values of $t$ and $\epsilon$.
$t$ $\epsilon$ $\frac{\|\widehat{X}_{CL}-X_{CL}\|_{F}}{\|X_{CL}\|_{F}}$ $U_{CL}$
$1$ $7.0103\times 10^{-13}$ $1.3055\times 10^{-12}$ $2.7846\times 10^{-11}$
$7.0580\times 10^{-10}$ $1.3349\times 10^{-9}$ $2.8036\times 10^{-8}$
$8.9885\times 10^{-7}$ $1.81513\times 10^{-6}$ $3.5704\times 10^{-6}$
$5$ $7.1245\times 10^{-11}$ $1.1503\times 10^{-10}$ $9.5002\times 10^{-9}$
$8.9399\times 10^{-9}$ $2.0884\times 10^{-8}$ $1.1921\times 10^{-6}$
$1.0173\times 10^{-8}$ $1.5357\times 10^{-8}$ $1.3565\times 10^{-6}$
$9$ $7.1497\times 10^{-12}$ $1.2430\times 10^{-11}$ $1.3370\times 10^{-9}$
$6.6856\times 10^{-10}$ $1.2308\times 10^{-9}$ $1.2502\times 10^{-7}$
$1.0212\times 10^{-7}$ $2.5587\times 10^{-7}$ $1.9096\times 10^{-5}$
Table 2: Comparison of relative forward errors and their upper bounds for the complex solution of a perturbed RBLSE problem (11)
Table 2 shows that the exact relative forward errors of the complex solution to the RBLSE problem (11) are consistently lower than their respective upper bounds across various values of $t$ and $\epsilon$. This verifies the reliability of the derived upper bound $U_{CL}$ for the relative forward error.
Example 5.3.
Let $A=A_{0}+A_{1}\textit{{i}}+A_{2}\textit{{j}}+A_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, $B=B_{0}+B_{1}\textit{{i}}+B_{2}\textit{{j}}+B_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $C=C_{0}+C_{1}\textit{{i}}+C_{2}\textit{{j}}+C_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n}$, and $D=D_{0}+D_{1}\textit{{i}}+D_{2}\textit{{j}}+D_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}$. Let
$A_{i}$ $=$ $randn(m,n)\in{\mathbb{R}}^{m\times n},\;\textrm{for}\;i=0:3,$
$B_{i}$ $=$ $randn(m,d)\in{\mathbb{R}}^{m\times d},\;\textrm{for}\;i=0:3,$
$C_{i}$ $=$ $randn(p,n)\in{\mathbb{R}}^{p\times n},\;\textrm{for}\;i=0:3,$
$D_{i}$ $=$ $randn(p,d)\in{\mathbb{R}}^{p\times d},\;\textrm{for}\;i=0:3.$
Take $m=30t$, $n=10t$, $p=2t$, and $d=2$. Here, $t$ is an arbitrary number. We apply Algorithm 2 to determine the real solution of the RBLSE problem. Let $X_{RL}$ be the real solution of the RBLSE problem
$\min_{X,R_{r}}\left\lVert R_{r}\right\rVert_{F}\;\;\textrm{subject to}\;\;AX=B+R_{r},\;CX=D.$
Let $\epsilon_{1}=\left\lVert AX_{RL}-(B+R_{r})\right\rVert_{F}$ and $\epsilon_{2}=\left\lVert CX_{RL}-D\right\rVert_{F}$. In Table 3, we compute $\epsilon_{1}$ and $\epsilon_{2}$ for different values of $t$.
$t$ $\epsilon_{1}$ $\epsilon_{2}$
$1$ $2.9063\times 10^{-15}$ $3.0851\times 10^{-15}$
$3$ $2.5053\times 10^{-15}$ $5.5184\times 10^{-15}$
$5$ $3.7915\times 10^{-15}$ $1.0949\times 10^{-14}$
$7$ $3.9204\times 10^{-15}$ $1.3185\times 10^{-14}$
$9$ $5.2026\times 10^{-15}$ $1.7247\times 10^{-14}$
Table 3: Computational accuracy of Algorithm 2 for computing the real solution of the RBLSE problem (29)
Table 3 shows that the errors $\epsilon_{1}$ and $\epsilon_{2}$ across different values of $t$ are consistently below $10^{-14}$. This indicates that Algorithm 2 is highly effective in determining the real solution for the RBLSE problem (29).
Example 5.4.
In Example 5.3, we introduce random perturbations $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ to the input matrices $A$, $B$, $C$, and $D$, respectively, to examine how these small perturbations affect the real solution $X_{RL}$ of the RBLSE problem (29). These perturbations are measured normwise by the smallest $\epsilon$, as in (38). We define the exact relative forward errors for these solutions as $\frac{\|\widehat{X}_{RL}-X_{RL}\|_{F}}{\|X_{RL}\|_{F}}$. Table 4 presents the computed exact relative forward errors and the corresponding upper bound $U_{RL}$ (calculated using Equation 40) for these solutions across different values of $t$ and $\epsilon$.
$t$ $\epsilon$ $\frac{\|\widehat{X}_{RL}-X_{RL}\|_{F}}{\|X_{RL}\|_{F}}$ $U_{RL}$
$1$ $7.3638\times 10^{-12}$ $2.3814\times 10^{-11}$ $8.0339\times 10^{-10}$
$7.3677\times 10^{-10}$ $2.1476\times 10^{-9}$ $8.0381\times 10^{-8}$
$7.9614\times 10^{-8}$ $7.6845\times 10^{-7}$ $8.6858\times 10^{-6}$
$3$ $7.1742\times 10^{-14}$ $2.6216\times 10^{-13}$ $9.0524\times 10^{-12}$
$8.1702\times 10^{-11}$ $1.5855\times 10^{-10}$ $1.0309\times 10^{-8}$
$7.1758\times 10^{-9}$ $2.6438\times 10^{-8}$ $9.0544\times 10^{-7}$
$5$ $7.2567\times 10^{-11}$ $2.4717\times 10^{-10}$ $1.0430\times 10^{-8}$
$7.2735\times 10^{-9}$ $2.5130\times 10^{-8}$ $1.0454\times 10^{-6}$
$8.7538\times 10^{-7}$ $1.8374\times 10^{-6}$ $1.2581\times 10^{-4}$
Table 4: Comparison of relative forward errors and their upper bounds for the real solution of a perturbed RBLSE problem (29)
Table 4 shows that the exact relative forward errors of the real solution to the RBLSE problem (29) are consistently lower than their respective upper bounds across various values of $t$ and $\epsilon$. This verifies the reliability of the derived upper bound $U_{RL}$ for the relative forward error.
Example 5.5.
Let $A=M_{1}+M_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times 10}$, $C=R_{1}+R_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{3\times 10}$, and $X_{0}=rand(10,3)+rand(10,3)\textit{{i}}\in{\mathbb{C}}^{10\times 3}$. Let
$M_{i}$ $=$ $rand(m,10)+rand(m,10)\textit{{i}}\in{\mathbb{C}}^{m\times 10},\;\textrm{for}\;i=1,2$
$R_{i}$ $=$ $rand(3,10)+rand(3,10)\textit{{i}}\in{\mathbb{C}}^{3\times 10},\;\textrm{for}\;i=1,2.$
Take $B=AX_{0}$ and $D=CX_{0}$. Thus, $X_{0}$ is the complex solution to the RBLSE problem (11). To measure the performance of the proposed technique for the RBLSE problem (11), Algorithm 1 is employed to determine the complex solution $X$. Let the error be $\epsilon=\left\lVert X-X_{0}\right\rVert_{F}$. Here, $m$ is a variable parameter. We evaluate the errors $\epsilon$ for various values of $m$. The relationship between the errors $\epsilon$ and $m$ is presented in Table 5.
$m$ $\epsilon=\|X-X_{0}\|_{F}$
$100$ $1.3154\times 10^{-14}$
$200$ $6.2150\times 10^{-15}$
$300$ $6.5603\times 10^{-15}$
$400$ $6.1485\times 10^{-15}$
$500$ $3.2441\times 10^{-15}$
Table 5: Computational accuracy of Algorithm 1 for computing the complex solution of the RBLSE problem (11)
Table 5 shows that the error $\epsilon$ between the complex solution derived from Algorithm 1 and the corresponding true solution to the RBLSE problem (11) remains consistently below $10^{-14}$ for various values of $m$. This indicates the high accuracy of Algorithm 1 in computing the complex solution to the RBLSE problem (11).
Example 5.6.
Let $A=A_{0}+A_{1}\textit{{i}}+A_{2}\textit{{j}}+A_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times 50}$, $C=C_{0}+C_{1}\textit{{i}}+C_{2}\textit{{j}}+C_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{10\times 50}$, and $X_{0}=randn(50,30)\in{\mathbb{R}}^{50\times 30}$. Let
$A_{i}$ $=$ $randn(m,50)\in{\mathbb{R}}^{m\times 50},\;\textrm{for}\;i=0:3$
$C_{i}$ $=$ $randn(10,50)\in{\mathbb{R}}^{10\times 50},\;\textrm{for}\;i=0:3.$
Take $B=AX_{0}$ and $D=CX_{0}$. Clearly, $X_{0}$ is the real solution to the RBLSE problem (29). To measure the performance of the proposed technique for the RBLSE problem (29), Algorithm 2 is employed to determine the real solution $X$. Define the error as $\epsilon=\left\lVert X-X_{0}\right\rVert_{F}$. Here, $m$ is a variable parameter. We evaluate the errors $\epsilon$ for various values of $m$. The relationship between the errors $\epsilon$ and $m$ is presented in Table 6.
$m$ $\epsilon=\|X-X_{0}\|_{F}$
$1000$ $3.8948\times 10^{-14}$
$2000$ $4.4732\times 10^{-14}$
$3000$ $4.1257\times 10^{-14}$
$4000$ $3.7531\times 10^{-14}$
$5000$ $4.6532\times 10^{-14}$
Table 6: Computational accuracies of Algorithm 2 for computing the real solution of the RBLSE problem (29)
Table 6 shows that the error $\epsilon$ between the real solution derived from Algorithm 2 and the corresponding true solution to the RBLSE problem (29) remains consistently below $10^{-14}$ for various values of $m$. This indicates the high accuracy of Algorithm 2 in computing the real solution to the RBLSE problem (29).
6 Conclusions
We have presented an algebraic method for solving the RBLSE problem by transforming it into complex and real LSE problems. This transformation leverages the complex and real representations of reduced biquaternion matrices, enabling efficient computation of both complex and real solutions to the RBLSE problem. Additionally, we have calculated the upper bound for the relative forward error associated with these solutions, demonstrating the robustness and accuracy of our method in addressing the RBLSE problem.
7 Declarations
Author’s Contributions
The first author played a crucial role in overseeing the project and providing valuable guidance. Neha Bhadala made significant contributions at various stages of the research, including conceptualization, methodology design, software implementation, validation, and formal analysis. Additionally, Neha Bhadala took the lead in drafting the original manuscript and actively participated in the subsequent writing and editing phases.
Acknowledgement
Neha Bhadala expresses gratitude to the Government of India for funding her research under the Prime Minister’s Research Fellowship program (PMRF ID: $210059$).